# Globalization of perturbative Chern-Simons theory on the moduli space of flat connections in the BV formalism

### v2

@@ 2.5.2. Desynchronized Hodge decomposition
Let $A_{0}\in\mathrm{FC}^{\mathrm{sm}}$ be a smooth flat connection. Then there is a neighborhood $U$ of $A_{0}$ in $\mathrm{FC}^{\mathrm{sm}}$ such that, for any $A^{\prime}\in U$, $d^{*}_{A^{\prime}}$ is a good gauge fixing operator for $d_{A_{0}}$.
@@
Let $A_{0}$ be a smooth flat connection and $A_{t}=A_{0}+\sum_{k\geq 1}t^{k}\alpha^{(k)}$ a path of smooth flat connections starting at $A_{0}$. Dually to Proposition 2.15, one can deform a harmonic form $\chi$ satisfying $d_{A_{0}}\chi=d^{*}_{A_{0}}\chi=0$ to a path $\widetilde{\chi}_{t}$ satisfying $d_{A_{0}}\widetilde{\chi}_{t}=d_{A_{t}}^{*}\widetilde{\chi}_{t}=0$, with
$\widetilde{\chi}^{(n)}=-d_{A_{0}}G_{A_{0}}\sum_{k+k^{\prime}=n,k\geq 1}\mathrm{ad}^{*}_{\alpha^{(k)}}\chi^{(k^{\prime})}.$
@@
    $W$ has constant (graded) rank and, since $W$ is contained in $\ker\mathcal{H}$ whose rank is non-increasing when moving $A^{\prime}$ away from $A_{0}$ (in an open neighborhood) and since $W=\ker\mathcal{H}$ at $A^{\prime}=A_{0}$, the rank of $\ker\mathcal{H}$ must stay constant. Hence, $\ker\mathcal{H}=W$ for $A^{\prime}\in U$.
@@
    $W$ contains a single representative of each $d_{A_{0}}$-cohomology class. Indeed, the map $q\colon W\rightarrow H_{A_{0}}$ sending $\alpha\mapsto[\alpha]$ is surjective, since Remark 2.29 defines a right inverse for $q$ – a map $\rho\colon H_{A_{0}}\rightarrow W$ satisfying $q\circ\rho=\mathrm{id}_{H_{A_{0}}}$. On the other hand, by (a) $W$ has constant rank when $A^{\prime}$ is changing, equal to the rank of $H_{A_{0}}$ at $A^{\prime}=A_{0}$. Hence, the fact that $q$ is a surjection implies that it is in fact an isomorphism.
@@ 3. Perturbative Chern-Simons partition function in the BV formalism
(93) $d_{A_{0}}=d+[A_{0},\cdot]\colon\Omega^{\bullet}(M,\mathfrak{g})\to\Omega^{\bullet+1}(M,\mathfrak{g})$
and the $A_{0}$-twisted de Rham cohomology by $H^{\bullet}_{d_{A_{0}}}(M,\operatorname{Ad}P)$.
@@ 4.1. Gauge invariance
– the r.h.s. of (115), as claimed.
@@ 4.3. Changing the gauge-fixing operator
(124) $Z_{A,A_{1}^{\prime}}(\mathsf{a})=Z_{A,A^{\prime}_{0}}(\mathsf{a})+i\hbar\Delta_{\mathsf{a}}R_{A,A^{\prime}_{0},A^{\prime}_{1}}(\mathsf{a}).$
@@
(127) $\delta_{A^{\prime}}Z_{A,A^{\prime}}(\mathsf{a})=i\hbar\Delta_{\mathsf{a}}\left(r_{A,A^{\prime};\delta A^{\prime}}(\mathsf{a})Z_{A,A^{\prime}}(\mathsf{a})\right)$
@@
Sketch of proof of Proposition 4.11.
We have
(129) $Z_{A,A^{\prime}}(\mathsf{a})^{-1}\,\delta_{A^{\prime}}Z_{A,A^{\prime}}(\mathsf{a})=\sum_{\Gamma\;\mathrm{connected}}\delta_{A^{\prime}}\overline{\Phi}_{\Gamma,A,A^{\prime}}(\mathsf{a})$
where we denoted
@@
$\overline{\Phi}_{\Gamma,A,A^{\prime}}(\mathsf{a})=\frac{(-i\hbar)^{-\chi(\Gamma)}}{|\mathrm{Aut}(\Gamma)|}\Phi_{\Gamma,A,A^{\prime}}(\mathsf{a})=\int_{\overline{\mathrm{Conf}}_{V}(M)}\omega_{\Gamma}$
@@
where $\omega_{\Gamma}$ is the integrand of (95) in $(A,A^{\prime})$-gauge, normalized with an appropriate power of $\hbar$ and a combinatorial factor.
The rest of the proof follows the arguments of [CM08], [CMR17].
Variation of the value of a Feynman graph in $A^{\prime}$ is the sum over edges and leaves of the graph of replacing that leaf with its variation, which contains terms $d,\Lambda]$, $d\mathbb{I}i$, see (126). Then one uses Stokes’ theorem on the configuration space to move the differential $d$ from the marked edge or leaf to other graph edges. The result is: $\delta_{A^{\prime}}\overline{\Phi}_{\Gamma}$ is a sum of
  1. (i)
    Graphs $\Gamma^{(1)}$ obtained by decorating one edge of $\Gamma$ with $\Lambda$ and one other edge by [d,K]=1P[d,K]=1-P.
  2. (ii)
    Graphs $\Gamma^{(2)}$ obtained by decorating one leaf of $\Gamma$ with $\mathbb{I}i(\mathsf{a})$ and one edge by 1P1-P.
  3. (iii)
    Graphs $\Gamma^{(3)}$ obtained by decorating one edge of $\Gamma$ with $\mathbb{I}P$.
Here we omit the subscripts in $d_{A},K_{A,A^{\prime}},P_{A,A^{\prime}}$. The contribution of $1$ on an edge can be seen as an integral over the principal boundary stratum of the configuration space (arising when we use the Stokes’ theorem as above) corresponding to a collapse of two points.
Figure 7. Examples of $\Gamma^{(1)}$, $\Gamma^{(2)}$, $\Gamma^{(3)}$. (We don’t write the contribution 11 on dashed edges as it cancels in the sum over graphs.)
@@
The contributions of $1$ on edges cancel out in the sum over graphs as a consequence of Jacobi identity (or, equivalently, as a consequence of the classical master equation on BV Chern-Simons action).
@@
Hidden boundary strata of the configuration space do not contribute by the standard arguments, see [CM08].
This implies that
@@
(130) $\sum_{\Gamma}\delta_{A^{\prime}}\overline{\Phi}_{\Gamma}=i\hbar\left(\Delta_{\mathsf{a}}\Big(\sum_{\Gamma^{\prime}}\overline{\Phi}_{\Gamma^{\prime}}\Big)+\Big\{\sum_{\Gamma^{\#}}\overline{\Phi}_{\Gamma^{\#}},\sum_{\Gamma^{\prime}}\overline{\Phi}_{\Gamma^{\prime}}\Big\}\right).$
Here $\Gamma$ and $\Gamma^{\#}$ runs over connected graphs. $\Gamma^{\prime}$ runs over connected grants with either (a) one edge decorated by $\Lambda$ or (b) one leaf decorated by $\mathbb{I}i(\mathsf{a})$.
Indeed, the $\delta_{A^{\prime}}\overline{\Phi}_{\Gamma}$ is the sum of decorated graphs $\Gamma^{(i)}$ as above, $i=1,2,3$ (we can ignore the contributions of 1 on edges, as discussed above). Removing the $P$-edge either
  1. (A)
    disconnects $\Gamma^{(i)}$ into a graph $\Gamma^{\prime}$ containing $\Lambda$ or $\mathbb{I}$ and a graph $\Gamma^{\#}$, or
  2. (B)
    does not disconnect the graph, and gives a connected graph $\Gamma^{\prime}$ containing $\Lambda$ or $\mathbb{I}$.
Adding back the $P$-edge (in all possible ways) amounts to applying $i\hbar\Delta_{\mathsf{a}}$ to $\overline{\Phi}_{\Gamma^{\prime}}$ in the case (B) and amounts to evaluating the Poisson bracket $i\hbar\{\overline{\Phi}_{\Gamma^{\#}},\overline{\Phi}_{\Gamma^{\prime}}\}$ in the case (A).
The factor $i\hbar=(-1)(-i\hbar)$ in the r.h.s. of (130) stems from the fact that Feynman weights contain the factor $(-i\hbar)^{-\chi(\Gamma)}$ and from the inclusion-exclusion formula for the Euler characteristic: adding back the $P$-edge changes $\chi$ by +1+1. The extra minus comes from the Stokes’ theorem on the configuration space.
From (130), denoting $\sum_{\Gamma^{\prime}}\overline{\Phi}_{\Gamma^{\prime}}$ by $r(\mathsf{a})$, we obtain
$Z_{A,A^{\prime}}(\mathsf{a})^{-1}\,\delta_{A^{\prime}}Z_{A,A^{\prime}}(\mathsf{a})=i\hbar\left(\Delta_{\mathsf{a}}r(\mathsf{a})+\Big\{\sum_{\Gamma^{\#}}\overline{\Phi}_{\Gamma^{\#}},r(\mathsf{a})\Big\}\right).$
Multiplying both sides by $Z_{A,A^{\prime}}(\mathsf{a})$, we obtain (127). ∎

### v3

@@ 2.5.2. Desynchronized Hodge decomposition
Let $A_{0}\in\mathrm{FC}^{\prime}$ be a smooth irreducible flat connection. Then there is a neighborhood $U$ of $A_{0}$ in $\mathrm{FC}^{\prime}$ such that, for any $A^{\prime}\in U$, $d^{*}_{A^{\prime}}$ is a good gauge fixing operator for $d_{A_{0}}$.
@@
Let $A_{0}$ be a smooth irreducible flat connection and $A^{\prime}_{t}=A_{0}+\alpha_{t}$, with $\alpha_{t}=\sum_{k\geq 1}t^{k}\alpha^{(k)}$, a path of flat connections starting at $A_{0}$. Dually to Proposition 2.23, one can deform a harmonic form $\chi$ satisfying $d_{A_{0}}\chi=d^{*}_{A_{0}}\chi=0$ to a path $\chi_{t}=\sum_{k\geq 0}t^{k}\chi^{(k)}$ with $\chi^{(0)}=\chi$, satisfying $d_{A_{0}}\chi_{t}=d_{A^{\prime}_{t}}^{*}\chi_{t}=0$, with
$\chi^{(n)}=-d_{A_{0}}G_{A_{0}}\sum_{k+k^{\prime}=n,k\geq 1}\mathrm{ad}^{*}_{\alpha^{(k)}}\chi^{(k^{\prime})}.$
Equivalently,
$\chi_{t}=\sum_{k\geq 0}(-d_{A_{0}}G_{A_{0}}\mathrm{ad}^{*}_{\alpha_{t}})^{k}\circ\chi.$
This statement follows by applying Hodge star to (70).
@@
    $W$ coincides with $\ker\mathcal{H}$ and has constant (graded) rank as $A^{\prime}\in U$ moves away from $A_{0}$. Indeed, one has:
    1. (i)
      $W\subset\ker\mathcal{H}$.
    2. (ii)
      At $A^{\prime}=A_{0}$, $W=\ker\mathcal{H}$ (by usual Hodge decomposition).
    3. (iii)
      Rank of $W$ is non-decreasing when moving $A^{\prime}$ away from $A_{0}$. Indeed, by Remark 2.34, one can deform a basis of $A_{0}$-harmonic forms to a linear independent set in $W$. (Since linear independence is an open condition.)
    4. (iv)
      Rank of $\ker\mathcal{H}$ is non-increasing as $A^{\prime}$ is moved away from $A_{0}$. – The rank of the kernel in a family of elliptic operator jumps (increases) on a closed subset in the space of parameters.
    These properties imply the claim above: $W=\ker\mathcal{H}$ and the rank is constant.
@@
    $W$ contains a single representative of each $d_{A_{0}}$-cohomology class. Indeed, the map $q\colon W\rightarrow H_{A_{0}}$ sending $\alpha\mapsto[\alpha]$ is surjective, since Remark 2.34 constructs a preimage under $q$ of a cohomology class $\alpha]$ as a deformation of its harmonic representative $\chi$ along a path $A^{\prime}_{t}$ from $A_{0}$ to $A^{\prime}$, to a form $\widetilde{\chi}$ satisfying $d_{A_{0}}\widetilde{\chi}=d^{*}_{A^{\prime}}\widetilde{\chi}=0$. On the other hand, by (a) $W$ has constant rank when $A^{\prime}$ is changing, equal to the rank of $H_{A_{0}}$ at $A^{\prime}=A_{0}$. Hence, the fact that $q$ is a surjection implies that it is in fact an isomorphism.
@@
Remark 2.36.
Notice that Proposition 2.31 is true also for the desynchronized Hodge gauge fixing, since we are only perturbing $\Delta_{A_{0}}$ by differential operators of degree one or less, so its inverse is still a degree -2 operator.
@@ 3. Perturbative Chern-Simons partition function in the BV formalism
(97) $d_{A_{0}}=d+[A_{0},\cdot]\colon\Omega^{\bullet}\to\Omega^{\bullet+1}$
and the $A_{0}$-twisted de Rham cohomology by $H_{A_{0}}^{\bullet}$.
@@ 4.1. Gauge invariance
– the r.h.s. of (119), as claimed. [^47] Here $\{f,g\}=f\left\langle\frac{\overleftarrow{\partial}}{\partial\mathsf{a}^{1}},\frac{\overrightarrow{\partial}}{\partial\mathsf{a}^{2}}\right\rangle g-f\left\langle\frac{\overleftarrow{\partial}}{\partial\mathsf{a}^{2}},\frac{\overrightarrow{\partial}}{\partial\mathsf{a}^{1}}\right\rangle g$ is the BV bracket of $f$ and $g$, see e.g. [Mne19, Section 4.4.3].
@@ 4.3. Changing the gauge-fixing operator
(128) $Z_{A,A_{1}^{\prime}}(\mathsf{a})=Z_{A,A^{\prime}_{0}}(\mathsf{a})-i\hbar\Delta_{\mathsf{a}}R_{A,A^{\prime}_{0},A^{\prime}_{1}}(\mathsf{a}).$
@@
(131) $\delta_{A^{\prime}}Z_{A,A^{\prime}}(\mathsf{a})=-i\hbar\Delta_{\mathsf{a}}\left(r_{A,A^{\prime};\delta A^{\prime}}(\mathsf{a})Z_{A,A^{\prime}}(\mathsf{a})\right)$
@@
Proof of Proposition 4.11.
Denote $Z^{0}_{A}=e^{\frac{i}{\hbar}S_{CS}(A)}e^{\frac{\pi i}{4}\psi_{A}}\tau_{A}^{\frac{1}{2}}$. We have
(133) $(Z^{0}_{A})^{-1}\delta_{A^{\prime}}Z_{A,A^{\prime}}(\mathsf{a})=\sum_{\Gamma}\delta_{A^{\prime}}\overline{\Phi}_{\Gamma,A,A^{\prime}}(\mathsf{a})$
– a sum over possibly disconnected trivalent graphs with leaves allowed, where we denoted
@@
$\overline{\Phi}_{\Gamma,A,A^{\prime}}(\mathsf{a})=\underbrace{\frac{(-i\hbar)^{-\chi(\Gamma)}}{|\mathrm{Aut}(\Gamma)|}}_{c_{\Gamma}}\Phi_{\Gamma,A,A^{\prime}}(\mathsf{a})=\int_{\overline{\mathrm{Conf}}_{V}(M)}\omega_{\Gamma}.$
@@
Here $\omega_{\Gamma}$ is the integrand of (99) in $(A,A^{\prime})$-gauge, normalized with an appropriate power of $\hbar$ and a combinatorial factor:
(134) $\omega_{\Gamma}=c_{\Gamma}\Big\langle\bigwedge_{e=(uv)\in E}\pi^{*}_{uv}\eta\wedge\bigwedge_{l\in L}\pi^{*}_{v(l)}i(\mathsf{a}),\bigotimes_{v\in V}f\Big\rangle.$
Here:
  • $V,E,L$ are the sets of vertices, edges and leaves of $\Gamma$.
  • For a short loop $e=(vv)$, expression $\pi_{vv}^{*}\eta$ is interpreted as $\pi_{v}^{*}\eta^{\Delta}$ with $\eta^{\Delta}$ the regularized diagonal evaluation of $\eta$.
  • $A,A^{\prime}$ subscripts are suppressed.
We will also use shorthand notation $\eta_{e}=\pi^{*}_{uv}\eta$, $i(\mathsf{a})_{l}=\pi_{v(l)}^{*}i(\mathsf{a})$.
The rest of the proof is parallel to the arguments of [CM08, Section 4.5, Appendix A], [CMR17, Lemma 4.11]. [^48] The technology of proving the properties of perturbative partition functions via Stokes’ theorem on the configuration space was developed in particular in [AS94, §6], [Kon94], [Kon03].
Variation of the value of a Feynman graph in $A^{\prime}$ is the sum over edges and leaves of the graph of replacing that edge or leaf with its variation:
(135) $\delta_{A^{\prime}}\overline{\Phi}_{\Gamma,A,A^{\prime}}=\\ \sum_{e_{\mathrm{m}}\in E}\int_{\overline{\mathrm{Conf}}_{V}(M)}c_{\Gamma}\Big\langle\delta_{A^{\prime}}\eta_{e_{\mathrm{m}}}\wedge\bigwedge_{e\in E\setminus e_{\mathrm{m}}}\eta_{e}\wedge\bigwedge_{l\in L}i(\mathsf{a})_{l},\bigotimes_{v\in V}f\Big\rangle\\ +\sum_{l_{\mathrm{m}}\in L}\int_{\overline{\mathrm{Conf}}_{V}(M)}c_{\Gamma}\Big\langle\bigwedge_{e\in E}\eta_{e}\wedge\delta_{A^{\prime}}i(\mathsf{a})_{l_{\mathrm{m}}}\wedge\bigwedge_{l\in L\setminus l_{\mathrm{m}}}i(\mathsf{a})_{l},\bigotimes_{v\in V}f\Big\rangle.$
(Subscript $\mathrm{m}$ in $e_{\mathrm{m}},l_{\mathrm{m}}$ is for “marked” edge/leaf.) We introduce the following notation: for $O\in\{\Lambda,P,\mathbb{I}P,P\mathbb{P}\}$ an operator on $\Omega^{\bullet}(M,\mathfrak{g})$, we will denote $\eta_{O}\in\Omega^{p}(\overline{\mathrm{Conf}}_{2}(M),\mathfrak{g}\otimes\mathfrak{g})$ its integral kernel, seen as a form on the compactified configuration space of two points of degree $p=1$ for $\Lambda$, $p=2$ for $\mathbb{I}P$ and $P\mathbb{P}$, $p=3$ for $P$. From (130) we have
(136) $\delta_{A^{\prime}}\eta=-d\eta_{\Lambda}+\eta_{\mathbb{I}P}+\eta_{P\mathbb{P}},\quad\delta_{A^{\prime}}i(\mathsf{a})=-d\mathbb{I}i(\mathsf{a}).$
Note that operators $\mathbb{I}P$ and $P\mathbb{P}$ are mutually transpose.
Next, we substitute (136) in (135) and use Stokes’ theorem on the configuration space to move $d$ from the marked edge $e_{\mathrm{m}}$ or leaf $l_{\mathrm{m}}$ to another edge $e^{\prime}_{\mathrm{m}}$ (note that [^49] As a distributional form on $M\times M$, the integral kernel of the operator [d,K]=1P[d,K]=1-P contains the integral kernel of the identity – the delta-distribution on the diagonal in $M\times M$. On the configuration space of pairs of distinct points on $M$, $M\times M\setminus\mathrm{Diag}$}, this delta-distribution disappears. However, in the formalism of configuration space integrals, it effectively reappears as the contribution of the principal boundary stratum of the configuration space, see (140), (141) and the discussion below it. $d\eta=\eta_{P}$, since [d,K]=1P[d,K]=1-P). In the process, Stokes’ theorem produces additional contributions of the boundary of the configuration space.
$\delta_{A^{\prime}}\overline{\Phi}_{\Gamma,A,A^{\prime}}=$
(137) $\sum_{e_{\mathrm{m}}\neq e^{\prime}_{\mathrm{m}}\in E}\int_{\overline{\mathrm{Conf}}_{V}(M)}c_{\Gamma}\cdot$
$\cdot\Big\langle(\eta_{P})_{e^{\prime}_{\mathrm{m}}}\wedge(\eta_{\Lambda})_{e_{\mathrm{m}}}\wedge\bigwedge_{e\in E\setminus\{e_{\mathrm{m}},e^{\prime}_{\mathrm{m}}\}}\eta_{e}\wedge\bigwedge_{l\in L}i(\mathsf{a})_{l},\bigotimes_{v\in V}f\Big\rangle$
(138) $+\sum_{l_{\mathrm{m}}\in L,\,e^{\prime}_{\mathrm{m}}\in E}\int_{\overline{\mathrm{Conf}}_{V}(M)}c_{\Gamma}\cdot$
$\cdot\Big\langle(\eta_{P})_{e^{\prime}_{\mathrm{m}}}\wedge\bigwedge_{e\in E\setminus e^{\prime}_{\mathrm{m}}}\eta_{e}\wedge\mathbb{I}i(\mathsf{a})_{l_{\mathrm{m}}}\wedge\bigwedge_{l\in L\setminus l_{\mathrm{m}}}i(\mathsf{a})_{l},\bigotimes_{v\in V}f\Big\rangle$
(139) $+\sum_{e_{\mathrm{m}}\in E}\int_{\overline{\mathrm{Conf}}_{V}(M)}c_{\Gamma}\cdot$
$\cdot\Big\langle(\eta_{\mathbb{I}P}+\eta_{P\mathbb{P}})_{e_{\mathrm{m}}}\wedge\bigwedge_{e\in E\setminus e_{\mathrm{m}}}\eta_{e}\wedge\bigwedge_{l\in L}i(\mathsf{a})_{l},\bigotimes_{v\in V}f\Big\rangle$
(140) $+\sum_{e_{\mathrm{m}}\in E}\int_{\partial\overline{\mathrm{Conf}}_{V}(M)}c_{\Gamma}\cdot$
$\cdot\Big\langle(\eta_{\Lambda})_{e_{\mathrm{m}}}\wedge\bigwedge_{e\in E\setminus e_{\mathrm{m}}}\eta_{e}\wedge\bigwedge_{l\in L}i(\mathsf{a})_{l},\bigotimes_{v\in V}f\Big\rangle$
(141) $+\sum_{l_{\mathrm{m}}\in L}\int_{\partial\overline{\mathrm{Conf}}_{V}(M)}c_{\Gamma}\cdot$
$\cdot\Big\langle\bigwedge_{e\in E}\eta_{e}\wedge\mathbb{I}i(\mathsf{a})_{l_{\mathrm{m}}}\wedge\bigwedge_{l\in L\setminus l_{\mathrm{m}}}i(\mathsf{a})_{l},\bigotimes_{v\in V}f\Big\rangle.$
@@
The boundary contributions here (140), (141) cancel out in the sum over graphs $\Gamma$ as follows. The boundary of the configuration space is stratified (see [AS94], [CM08]) into “principal strata” (differential geometric blow-ups of collapses of two points on $M$) and “hidden strata” (collapses of $\geq 3$ points). Contributions of principal strata cancel out in the sum over graphs $\Gamma$ as a consequence of Jacobi identity (or, equivalently, as a consequence of the classical master equation on BV Chern-Simons action).
@@
Hidden boundary strata of the configuration space do not contribute to the integrals in (140), (141); we defer the proof of this to Appendix D.3.
Figure 7. Examples of graphs contributing to (137), (138), (139).
Denote
(142) $\zeta=\\ \sum_{\Gamma^{\#}}\sum_{e_{\mathrm{m}}\in E^{\#}}\int_{\overline{\mathrm{Conf}}_{V^{\#}}(M)}c_{\Gamma^{\#}}\Big\langle(\eta_{\Lambda})_{e_{\mathrm{m}}}\wedge\bigwedge_{e\in E^{\#}\setminus e_{\mathrm{m}}}\eta_{e}\wedge\bigwedge_{l\in L^{\#}}i(\mathsf{a})_{l},\bigotimes_{v\in V^{\#}}f\Big\rangle\\ +\sum_{\Gamma^{\#}}\sum_{l_{\mathrm{m}}\in L^{\#}}\int_{\overline{\mathrm{Conf}}_{V^{\#}}(M)}c_{\Gamma^{\#}}\Big\langle\bigwedge_{e\in E^{\#}}\eta_{e}\wedge\mathbb{I}i(\mathsf{a})_{l_{\mathrm{m}}}\wedge\bigwedge_{l\in L^{\#}\setminus l_{\mathrm{m}}}i(\mathsf{a})_{l},\bigotimes_{v\in V^{\#}}f\Big\rangle.$
Here $\Gamma^{\#}$ runs over possibly disconnected trivalent graphs with leaves allowed; $V^{\#},E^{\#},L^{\#}$ stand for the sets of vertices/edges/leaves of $\Gamma^{\#}$.
Note that
(143) $\zeta=(Z^{0}_{A})^{-1}r_{A,A^{\prime};\delta A^{\prime}}(\mathsf{a})Z_{A,A^{\prime}}(\mathsf{a})$
with $r$ defined as in Proposition 4.11.
On the other hand, applying $\Delta_{\mathsf{a}}$ to each term in (142) yields a sum over pairs of $i(\mathsf{a})$ leaves, of replacement of that pair by a new edge adjoined to the graph $\Gamma^{\#}$ and decorated by $P$:
$\Delta_{\mathsf{a}}\zeta=$
(144) $\sum_{\Gamma^{\#}}\sum_{e_{\mathrm{m}}\in E^{\#},\,l_{\mathrm{m}}\neq l^{\prime}_{\mathrm{m}}\in L^{\#}}\int_{\overline{\mathrm{Conf}}_{V^{\#}}(M)}c_{\Gamma^{\#}}\cdot$
$\cdot\Big\langle(\eta_{P})_{(v(l_{\mathrm{m}})v(l^{\prime}_{\mathrm{m}}))}\wedge(\eta_{\Lambda})_{e_{\mathrm{m}}}\wedge\bigwedge_{e\in E^{\#}\setminus e_{\mathrm{m}}}\eta_{e}\wedge\bigwedge_{l\in L^{\#}\setminus\{l_{\mathrm{m}},l^{\prime}_{\mathrm{m}}\}}i(\mathsf{a})_{l},\bigotimes_{v\in V^{\#}}f\Big\rangle$
(145) $+\sum_{\Gamma^{\#}}\sum_{l_{\mathrm{m}},l^{\prime}_{\mathrm{m}},l^{\prime\prime}_{\mathrm{m}}\in L^{\#}\,\mathrm{distinct}}\int_{\overline{\mathrm{Conf}}_{V^{\#}}(M)}c_{\Gamma^{\#}}\cdot$
$\cdot\Big\langle\bigwedge_{e\in E^{\#}}\eta_{e}\wedge\mathbb{I}i(\mathsf{a})_{l_{\mathrm{m}}}\wedge(\eta_{P})_{(v(l^{\prime}_{\mathrm{m}})v(l^{\prime\prime}_{\mathrm{m}}))}\wedge\bigwedge_{l\in L^{\#}\setminus\{l_{\mathrm{m}},l^{\prime}_{\mathrm{m}},l^{\prime\prime}_{\mathrm{m}}\}}i(\mathsf{a})_{l},\bigotimes_{v\in V^{\#}}f\Big\rangle$
(146) $+\sum_{\Gamma^{\#}}\sum_{l_{\mathrm{m}}\neq l^{\prime}_{\mathrm{m}}\in L^{\#}}\int_{\overline{\mathrm{Conf}}_{V^{\#}}(M)}c_{\Gamma^{\#}}\cdot$
$\cdot\Big\langle\bigwedge_{e\in E^{\#}}\eta_{e}\wedge(\eta_{\mathbb{I}P})_{(v(l_{\mathrm{m}})v(l^{\prime}_{\mathrm{m}}))}\wedge\bigwedge_{l\in L^{\#}\setminus\{l_{\mathrm{m}},l^{\prime}_{\mathrm{m}}\}}i(\mathsf{a})_{l},\bigotimes_{v\in V^{\#}}f\Big\rangle.$
Comparing with the computation of $\delta_{A^{\prime}}\overline{\Phi}_{\Gamma,A,A^{\prime}}$, we find that:
[…]
Therefore we have
@@
(147) $\underbrace{\sum_{\Gamma}\delta_{A^{\prime}}\overline{\Phi}_{\Gamma,A,A^{\prime}}}_{(Z^{0}_{A})^{-1}\delta_{A^{\prime}}Z_{A,A^{\prime}}(\mathsf{a})}=\underbrace{-i\hbar\Delta_{\mathsf{a}}\zeta}_{i\hbar(Z^{0}_{A})^{-1}\Delta_{\mathsf{a}}(r_{A,A^{\prime};\delta A^{\prime}}(\mathsf{a})Z_{A,A^{\prime}}(\mathsf{a}))}$
Multiplying both sides by $Z^{0}_{A}$, we have (131).